OpenStudy (anonymous):

use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4

OpenStudy (anonymous):

you just replace y with 3x+4 in the first equation so instead of 2x+y=6, you have 2x+3x+4=6 so 5x=6-4 x=?

OpenStudy (anonymous):

thank you purple rainbow cherry